An airplane, flying against the wind dierction, travels 3000 miles between 2 cities in 6 hrs and 15 mts. In the return flight it travels the same distance but in 5 hrs.
Find the velocity of the plane in still air and the wind velocity, assuming that both remained constant in the round trip.

Answer by checkley71(8403)

3000/6.25=480 PLANE SPEED AGAINST THE WIND
3000/5=600 PLANE SPEED WITH THE WIND
(600-480)/2=120/2=60 MPH FOR THE WIND
PROOF:
600-60=480+60
540=540 SPEED FO THE PLANE IN STILL AIR.